1. 2,a,b,9. The first 3 terms form an Arithmetic Progression and the last 3 terms form form a Geometric Progression. Determine the values of a and b.
1
Expert's answer
2020-07-01T07:09:55-0400

In an Arithmetic progression ;the first term is a,the second a+d, the third a+2d and the fourth a+3d.

In this case a=2 and a+3d =9

If you replace:2+3d=9.substituting d=2.33.

Therefore the second term which in this case is a,will be given by a+d.Which is 2+2.33=4.33

The third term b;a+2d which is 2+4.66=6.66

Hence,a=4.44 and b= 6.66